Two cars start from points 400 miles apart and travel toward each other. They meet after 4 hours. Find the average speed of each car if one travels 20 miles per hour faster than the other.
Answer by mananth(16949)

Car I x mph
Car II x+ 20
..
They are moving towards each other

So add up their speed.

combined speed = x+ 1 x + 20
( 2 x + 20 )
Time = 4 hours
Distance = 400 miles
Distance = speed * time
( 2 x 20 )* 4 = 400
8 x + 80 = 400
8 x = 400 + -80
8 x = 320
/ 8
x= 40 mph Car I
Car II 40 + 20 = 60 mph
m.ananth@hotmail.ca